OpenStudy (lgbasallote):

\[\int_0^2 \frac{dx}{\sqrt{2x - x^2}}\] i got a 6 out of 12 points here i dont know why =_= btw my answer is \(2\pi\)

myininaya (myininaya):

\[2x-x^2=-x^2+2x=-(x^2-2x)=-(x^2-2x+1)+1 =-(x-1)^2+1\] Was this the way you wrote your bottom inside that square root thingy?

OpenStudy (lgbasallote):

uhmm yep \(1 - (x-1)^2\)

OpenStudy (lgbasallote):

\[\large \lim_{t \rightarrow 0} \int_t^1 \frac{dx}{\sqrt{1 - (x-1)^2}}\]

OpenStudy (auctoratrox):

now just do a trig sub and you got it

OpenStudy (lgbasallote):

\[\large \lim_{t \rightarrow 2} \int_1^t \frac{dx}{\sqrt{1 - (x-1)^2}}\]

OpenStudy (lgbasallote):

those are my limits

OpenStudy (lgbasallote):

where did i go wrong @myininaya

myininaya (myininaya):

\[\int\limits_{}^{}\frac{dx}{\sqrt{1-(x-1)^2}}\]  I chose the labeling that way so that when I find the other side it comes out to be the bottom :)  \[\text{ Let } \cos(\theta)=\frac{x-1}{1} (=\frac{adj}{hyp})\] \[\cos(\theta) =x-1 => -\sin(\theta) d \theta =dx\] \[\int\limits_{}^{}\frac{-\sin(\theta) d \theta}{\sqrt{1-\cos^2(\theta)}}=\int\limits_{}^{}-1 d \theta=-\theta+C\] \[=-\cos^{-1}(x-1)+C\] So we need to look back at the limits right ?